OpenStudy (anonymous):

What is the value of f[g(6)] for the functions f(x) = 2x + 1 and g(x) = 2x - 5?

OpenStudy (anonymous):

f(x)=2x+1_g(x)=2x-6_f(g(1)) Multiply 2 by each term inside the parentheses. f(2x-6)=4x-12+1 Add 1 to -12 to get -11. f(2x-6)=4x-11 Simplify the evaluated function. -7

OpenStudy (anonymous):

Oops, no wonder I was getting it wrong. I kept dividing them by 2... thanks!